6) A rock is thrown vertically with an initial velocity of 24.5 m s-1. What maximum height will the rock reach?



1
Expert's answer
2021-03-10T17:14:23-0500

Let's find the time that the rock takes to reach the maximum height:


"v_y=v_{0}-gt,""0=v_0-gt,""t=\dfrac{v_0}{g}."

We can find the maximum height reached by the rock from the kinematic equation:


"y_{max}=v_0t-\dfrac{1}{2}gt^2."

Substituting "t" into the formula, we get:


"y_{max}=v_0\dfrac{v_0}{g}-\dfrac{1}{2}g(\dfrac{v_0}{g})^2,""y_{max}=\dfrac{v_0^2}{2g}=\dfrac{(24.5\ \dfrac{m}{s})^2}{2\cdot9.8\ \dfrac{m}{s^2}}=30.62\ m."